# Coherence-Mediated Quantum Thermometry in a Hybrid Circuit-QED Architecture

### v1

@@ II.4 Strategy 2: Probe Phase-Shift Tracking
Building on the qubit-based coherence measurement in Strategy 1, the same temperature-dependent phase $\phi_{b}(T)=\lambda\tau\bar{n}_{a}(T)$ can be viewed as the parameter that would accumulate during a Ramsey sequence on a dispersively coupled qubit. Alternatively, we measure the phase variation directly via heterodyne detection of the high-$Q$ probe cavity, instead of mapping it onto a qubit superposition.
In this heterodyne-based implementation, the cross-Kerr interaction between the thermal mode and probe induces a temperature-dependent phase shift of the intracavity field. This phase is faithfully transferred to the outgoing field and measured through heterodyne detection of its quadratures. Because the amplitude remains unchanged, the probe visibility is unity,
$\mathcal{V}_{\Phi}(\tau,T)=1,$ (8)
and the information about temperature resides purely in the phase. The corresponding quantum Fisher information,
$\mathcal{F}_{\Phi}(T)=\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2},$ (9)
@@
This approach leverages the exceptional phase stability of ultra-high-$Q$ 3D cavities ($\kappa_{b}\!\ll\!$ kHz) to permit millisecond-scale integration times without qubit-coherence limitations. Thus, it naturally supports continuous, low-bandwidth monitoring of slow thermodynamic drifts with state-of-the-art sensitivity.
@@
Taken together, the phase-shift strategy complements the coherence-mediated approach. While Strategy 1 enables fast, high-bandwidth thermometry with short interaction times, Strategy 2 exploits long-lived cavity coherence for ultra-high sensitivity and naturally extends to calorimetry, where discrete energy events are registered as probe phase jumps.
@@ III Experimental Implementation and Feasibility
Building on the theoretical framework outlined above, we now describe a realistic experimental implementation of the coherence-mediated quantum thermometry scheme.
@@ III.1 Circuit Architecture and Mode Realization
The proposed thermometer comprises three essential elements: a high-coherence superconducting qubit, a low-frequency thermal mode, and a high-frequency coherent probe mode. These are integrated in a hybrid architecture combining planar CPW resonators with ultra-high-$Q$ 3D cavity components.
@@
The thermal channel is realized as a deliberately lossy CPW resonator or lumped LC circuit at 121{-}2 GHz. Operating at lower frequencies ensures a sizable Bose occupation $\bar{n}_{a}$ even at tens of millikelvin. By coupling strongly to an engineered bath, such as a terminated transmission line or resistor load, the mode rapidly equilibrates with its environment [24, 45], making it a faithful temperature sensor.
@@
In the coherence-mediated scheme, the qubit remains in its ground state while thermal fluctuations imprint dephasing noise onto the long-lived probe during an interaction interval $\tau$. A subsequent Ramsey-type sequence maps this probe dephasing back to the qubit for readout. In the phase-shift scheme, the probe field itself is monitored continuously via heterodyne detection; its ultra-high $Q$ enhances sensitivity to stochastic fluctuations in the thermal mode occupancy $\bar{n}_{a}(T)$.
@@ III.2.1 Engineering Enhanced Cross–Kerr
[…]
@@
In our scheme, the effective cross–Kerr must be strong enough ($\lambda/2\pi\sim 10$$50~\mathrm{kHz}$) to imprint temperature–dependent fluctuations onto the probe within a realistic interaction time $\tau$, while remaining compatible with achievable detunings and coherence in a hybrid circuit–QED architecture.
Conventional approaches use four–wave mixing to enhance the interaction between modes $a$ and $b$. An off–resonant pump can dress an existing transmon and increase the cross–Kerr without new hardware [31, 48]. Additionally, driven couplers placed between the modes can offer comparable, tunable $\lambda$ [49, 50, 51]. While flexible, these methods introduce AC–Stark and Kerr drifts and can raise the inverse–Purcell/noise floor seen by a high–$Q$ probe, adding calibration and packaging overhead that is difficult to reconcile with a 3D–cavity implementation.
@@
We therefore implement $\lambda\,\hat{n}_{a}\hat{n}_{b}$ using two fixed–frequency transmons (Q1, Q2) that passively bridge the thermal mode $a$ and the probe $b$ (see Fig. 1). Mode $a$ couples dispersively to Q1 with pull $\chi_{a1}=g_{a1}^{2}/\Delta_{a1}$, mode $b$ couples to Q2 with pull $\chi_{b2}=g_{b2}^{2}/\Delta_{b2}$, and the qubits are capacitively linked with an XY exchange $J_{XY}$. In the deep–dispersive regime, $|g_{a1,b2}|\ll|\Delta_{a1,b2}|,\quad|J_{XY}|\ll|\Delta_{12}|,$ the virtual coupling chain $a-Q_{1}-Q_{2}-b$ yields an effective cross–Kerr (Appendix E)
$\lambda\;=\;\frac{8\,\chi_{a1}\,\chi_{b2}\,J_{XY}^{2}}{\Delta_{12}^{3}}\;\times\;\big[1+\mathcal{O}(\tfrac{\Delta_{12}}{\Delta_{a1}},\tfrac{\Delta_{12}}{\Delta_{b2}})\big],$ (10)
where the order–unity corrections reflect finite transmon anharmonicities and detuning signs [31, 52]. Figure 4 shows the effective cross-Kerr $\lambda$ as a function of dispersive pulls $\chi_{a1,b2}$. For a fixed dispersive ratio $\rho\equiv J_{XY}/\Delta_{12}$, the scaling rewrites as $\lambda=8\,\chi_{a1}\chi_{b2}\rho^{2}/\Delta_{12}$, so increasing $\chi_{a1}$ (thermal side) and/or reducing $\Delta_{12}$ most efficiently boosts $\lambda$ while keeping $\chi_{b2}$ modest to protect the probe $Q$.
[…]
III.2.2 Impact of Qubit–Mode Couplings on Visibility
Figure 5: Qubit Ramsey visibility as a function of dispersive coupling strengths during the Ramsey window $\tau_{R}$. (a) Dependence on $\chi_{b}/2\pi$ for 10, 20, 5010,\,20,\,50 kHz. (b) Dependence on $\chi_{a}/2\pi$ for 1, 5, 101,\,5,\,10 kHz. The y-axis shows the normalized Ramsey visibility (qubit contrast) versus the Ramsey integration time $\tau_{R}$. Larger dispersive couplings increase dephasing during readout, shortening the usable $\tau_{R}$ window.
In the coherence-mediated scheme, the qubit is prepared and held in its ground state throughout the imprinting interval $\tau$, so it does not acquire dephasing from either mode during sensing. Visibility becomes relevant only during the subsequent Ramsey readout of duration $\tau_{R}$, when the qubit is placed in a superposition to sample the probe’s phase fluctuations.
If the qubit couples dispersively to a bosonic mode $m\in\{a,b\}$ with photon-number variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$, a second-order cumulant expansion gives the coherence envelope during the Ramsey window (Appendix F)
$C_{q}^{(m)}(\tau_{R})=\exp\!\left\{-\,4\chi_{m}^{2}\,\mathrm{Var}(n_{m})\left[\frac{\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}}}{\kappa_{m}^{2}}\right]\right\},$
with $\mathrm{Var}(n_{b})=|\alpha|^{2}+\bar{n}_{b}(1+\bar{n}_{b})$ for the probe and $\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a})$ for the thermal mode. The coherent amplitude $|\alpha|^{2}$ contributes photon shot noise.
[…]
In summary, no dynamical decoupling is required during sensing because the qubit stays in $|g\rangle$ throughout $\tau$. The only exposure to dephasing occurs in the short Ramsey interval $\tau_{R}$, where a large $\chi_{b}$ enables fast, high-fidelity readout, while a small $\chi_{a}$ avoids parasitic dephasing.
@@ IV Comparative Performance of Quantum Thermometry Strategies
Figure 6: Comparison of thermometric strategies. (a) Quantum Fisher information (QFI) $\mathcal{F}$ as a function of interaction time $\tau$ for the coherence-mediated (blue), phase-shift (orange), and qubit-only Ramsey (green) schemes. (b) Normalized coherence envelopes. The blue curve shows the probe envelope $C(\tau,T)$ in the coherence-mediated scheme, which is mapped to the qubit through dispersive readout. The green curve shows the qubit’s direct Ramsey envelope under thermal dephasing. The phase-shift scheme maintains unit contrast (orange). Together, the plots highlight that the coherence-mediated scheme achieves higher QFI at short times, while visibility decay ultimately limits long-$\tau$ sensitivity.
To benchmark performance, we compare the QFI and the associated visibility across three strategies: coherence-mediated sensing, direct phase-shift tracking, and qubit-only Ramsey thermometry [7, 8]. Details of the qubit-only scheme are provided in Appendix D. Figure 6 summarizes the results.
Panel (a) shows the QFI $\mathcal{F}_{T}(\tau)$ at $T=10$ mK. The coherence-mediated strategy (blue) exhibits the largest QFI at short and intermediate times, owing to the probe’s ability to convert thermal number fluctuations into amplified phase diffusion. However, this gain collapses at long $\tau$ as the probe visibility decays. The phase-shift strategy (orange) accumulates information linearly with $\tau$ and avoids dephasing, yielding steadily increasing QFI but with smaller values at short times. The qubit-only scheme (green) lies in between: it achieves an intermediate maximum QFI that can surpass the phase-shift strategy at moderate $\tau$, but is ultimately limited by qubit dephasing from the thermal mode.
Panel (b) displays the corresponding visibilities. The coherence-mediated scheme suffers the fastest decay due to probe dephasing. The qubit-only scheme also loses visibility, though more gradually. By contrast, the phase-shift scheme maintains unit visibility at all times, since the signal resides entirely in deterministic phase accumulation.
Taken together, these results highlight complementary trade-offs among the three strategies. The coherence-mediated scheme uniquely exploits the qubit’s quantum coherence as a transient amplifier, transducing thermal fluctuations into enhanced phase sensitivity. Within the weak-dephasing window ($\tau\lesssim 10~\mu\mathrm{s}$ for the chosen parameters, extendable by reducing the cross-Kerr $\lambda$), this approach achieves the strongest instantaneous quantum Fisher information and the steepest signal slope, outperforming both qubit-only and phase-shift tracking schemes in the short-time regime. In contrast, the qubit-only method is fundamentally limited by intrinsic dephasing, leading to a comparable visibility–sensitivity trade-off at longer $\tau$. The phase-shift strategy is the most robust, supporting arbitrarily long interaction times, though it lacks the amplification benefit of a coherent probe. Thus, coherence-mediated thermometry provides the greatest quantum gain, contingent on a high-$Q$ probe and precise dephasing control, both of which are well within current experimental reach, making it a promising route for practical quantum thermometry.
The essential difference between these two schemes lies in how temperature information is stored. In the qubit-only approach, the qubit is both sensor and memory: the same thermal-induced dephasing that encodes $T$ simultaneously destroys qubit visibility as $\tau$ grows. Because this dephasing originates from the sensing channel itself, dynamical decoupling cannot mitigate it without erasing the signal. By contrast, in the coherence-mediated scheme, the qubit can remain isolated during the sensing interval while thermal fluctuations are mapped onto the probe. The qubit is only re-engaged briefly for readout, protected by dynamical decoupling or idling during interaction. A high-$Q$ probe cavity suppresses visibility loss, enabling longer interaction times before decoherence dominates. In this way, the coherence-mediated strategy shifts the fundamental limitation away from the qubit’s intrinsic coherence and into a regime where cavity quality factors and readout protocols can be engineered for improved sensitivity.
Beyond raw sensitivity, a practical figure of merit is the information gained per unit wall-clock time. We define the Fisher-information rate as $\dot{\mathcal{F}}_{T}(\tau)\equiv\mathcal{F}_{T}(\tau)/(\tau+\tau_{\mathrm{oh}}),$ where $\tau_{\mathrm{oh}}$ accounts for preparation and readout overhead.
For the coherence-mediated and qubit-only schemes, the scaling $\mathcal{F}_{T}\!\sim\!\tau^{2}e^{-2\Gamma\tau}$ yields an optimal interaction time $\tau^{\star}\!\sim\!1/(2\Gamma)$ that maximizes $\dot{\mathcal{F}}_{T}$ and defines an effective measurement bandwidth $B^{\star}\!\sim\!\Gamma$. In contrast, the phase-shift strategy, with $\mathcal{F}_{T}\!\sim\!\tau^{2}$ and no visibility penalty, supports a continuously increasing $\dot{\mathcal{F}}_{T}(\tau)$ up to the limit set by technical drifts. Rather than a bandwidth constraint, this reflects its ability to integrate temperature information coherently over long durations, making it ideally suited for quantum calorimetry where slow thermal relaxation or discrete energy-deposition events are to be resolved with maximal precision. Thus, the phase-tracking scheme complements the high-bandwidth coherence-mediated approach by providing a low-noise, high-fidelity pathway toward long-term thermodynamic monitoring.
This comparison shows that the coherence-mediated and qubit-only protocols naturally operate in a high-bandwidth regime, making them well suited for fast thermometry and real-time temperature tracking. By contrast, the phase-shift strategy excels in the opposite limit: it enables coherent accumulation of temperature information over long durations without visibility loss, providing the optimal platform for quantum calorimetry and other applications requiring ultra-stable, long-integration measurements.

### v2

@@ II.4 Strategy 2: Probe Phase-Shift (PS) Tracking
Building on the qubit-based coherence measurement in Strategy 1, the same temperature-dependent phase $\phi_{b}(T)=\lambda\tau\bar{n}_{a}(T)$ can be viewed as the parameter that would accumulate during a Ramsey sequence on a dispersively coupled qubit. Alternatively, we measure the phase variation directly via heterodyne (or homodyne) detection of the high-$Q$ probe cavity, instead of mapping it onto a qubit superposition.
In this heterodyne-based implementation, the cross-Kerr interaction between the thermal mode and probe induces a temperature-dependent phase shift of the intracavity field. This phase is faithfully transferred to the outgoing field and measured through heterodyne detection of its quadratures. Because the amplitude remains unchanged, the probe visibility is unity $\mathcal{V}_{\Phi}(\tau,T)=1,$ and the information about temperature resides purely in the phase. The corresponding QFI,
$\mathcal{F}_{\Phi}(T)=4\alpha^{2}\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2},$ (7)
@@
This approach leverages the exceptional phase stability of high-$Q$ 3D cavities ($\kappa_{b}\!\ll\!$ kHz) to permit millisecond-scale integration times without qubit-coherence limitations. Thus, it naturally supports continuous, low-bandwidth monitoring of slow thermodynamic drifts with state-of-the-art sensitivity.
@@
We therefore view the PS strategy as complementary to the CM approach. While Strategy 1 enables fast, high-bandwidth thermometry with short interaction times, Strategy 2 exploits long-lived cavity coherence for ultra-high sensitivity and naturally extends to calorimetry, where discrete energy events are registered as probe phase jumps.
@@ III Experimental Implementation and Feasibility
Building on the theoretical framework outlined above, we now describe a realistic experimental implementation of the CM quantum thermometry scheme. In addition to outlining a concrete device architecture, this section summarizes the main practical constraints at a coarse-grained level and how they enter the protocol. Importantly, these effects primarily reduce visibility and set an optimal interaction/mapping time, but do not alter the central operating principle of the scheme: a long-lived coherent reference mediates the readout of weak thermal fluctuations while the qubit remains largely protected during the sensing interval.
@@ III.1 Circuit Architecture and Mode Realization
The proposed thermometer comprises three essential elements: a high-coherence superconducting qubit, a low-frequency thermal mode, and a high-frequency coherent probe mode. These are integrated in a hybrid architecture combining planar CPW resonators with high-$Q$ 3D cavity components.
@@
The thermal channel is realized as a planar CPW resonator or lumped LC circuit at 121{-}2 GHz [69, 13]. Operating at lower frequencies ensures a sizable thermal population $\bar{n}_{a}$ even at tens of millikelvin. By coupling to an engineered bath, such as a terminated transmission line or resistor load, the mode rapidly equilibrates with its environment [28, 65], ensuring that it faithful tracks the bath temperature.
@@
In the CM scheme, the qubit remains in its ground state while thermal fluctuations imprint dephasing noise onto the long-lived probe during an interaction interval $\tau$. A subsequent Ramsey-type sequence maps this probe dephasing back to the qubit for readout (Appendix G). In the PS scheme, the probe field itself is monitored continuously via heterodyne detection; its ultra-high $Q$ enhances sensitivity to stochastic fluctuations in the thermal mode occupancy $\bar{n}_{a}(T)$. In practice, heterodyne detection efficiency $\eta<1$ reduces the observed Fisher information compared to the quantum limit, while homodyne measurement of a single quadrature can approach unity.
@@ III.2.1 Engineering Enhanced Cross–Kerr
Figure 4: Heatmap of the two-qubit–mediated cross–Kerr rate $\lambda/2\pi$ (kHz) versus the dispersive pulls $\chi_{a1}/2\pi$ and $\chi_{b2}/2\pi$ (MHz). Values are computed from $\lambda=8\,\chi_{a1}\chi_{b2}\rho^{2}/\Delta_{12}$ with $\rho=0.2$ and $\Delta_{12}/2\pi=100~\mathrm{MHz}$. White dashed contours indicate $\lambda/2\pi=\{10,20,50,100\}$ kHz.
@@
In the scheme, the effective cross–Kerr $\lambda$ can be engineered to imprint temperature–dependent fluctuations onto the probe within a realistic interaction time $\tau$, while remaining compatible with achievable detunings and coherence in a hybrid cQED architecture.
A single transmon dispersively coupled to both modes can generate a fourth-order virtual cross-Kerr interaction [4, 66]. For realistic couplings and detunings this effective rate remains in the sub-kilohertz range, far too small for our sensing requirements, while attempts to enhance it by increasing hybridization inevitably degrade the probe’s quality factor and introduce additional Purcell and AC-Stark channels.
@@
To overcome this limitation, we employ a passive two-qubit bridge (Q1, Q2) that links the thermal mode $a$ and the probe mode $b$ (Fig. 1). Mode $a$ couples dispersively to Q1 with pull $\chi_{a1}$; mode $b$ couples to Q2 with pull $\chi_{b2}$; and the qubits are capacitively linked through an exchange interaction $J_{XY}$. In the dispersive regime, $|g_{a1,b2}|\ll|\Delta_{a1,b2}|,\qquad|J_{XY}|\ll|\Delta_{12}|,$ virtual transitions through the chain $a\!\leftrightarrow\!Q_{1}\!\leftrightarrow\!Q_{2}\!\leftrightarrow\!b$ generate an effective cross-Kerr (Appendix E)
$\lambda=\frac{8\chi_{a1}\chi_{b2}J_{XY}^{2}}{\Delta_{12}^{3}}.$ (8)
Here, we ignore the subleading terms in Eq. (33) that account for finite anharmonicity and detuning sign [4, 21]. Introducing the dimensionless ratio $\rho=J_{XY}/\Delta_{12}$, the scaling reduces to $\lambda=8\,\chi_{a1}\chi_{b2}\rho^{2}/\Delta_{12}.$ Figure 4 shows that a moderate $\lambda$ in 10105050 kHz regime is easily reached by increasing $\chi_{a1}$ while maintaining modest $\chi_{b2}$ to protect the probe’s quality factor.
A central advantage of the passive two-qubit bridge is that it leaves the 3D cavity nearly unaffected. With the state-of-the-art transmon lifetime $T_{1}\geq 300~\mu s$ [3, 47, 70] and in the dispersive regime, the dominant added loss arises from Q2’s inverse-Purcell contribution, $\kappa_{b,\mathrm{add}}/2\pi\approx(g_{b2}/\Delta_{b2})^{2}/(2\pi T_{1})\ll 1~\mathrm{kHz}$, far below the intrinsic linewidth of a high-$Q$ cavity. Unwanted $a\!\leftrightarrow\!b$ exchange is suppressed by ordering the qubit frequencies ($\omega_{1}<\omega_{2}$), leading to destructive interference between opposite-sign virtual paths. Meanwhile, the desired $\lambda$ remains large because it scales as $J_{XY}^{2}$.
For completeness, we note that alternative routes, such as parametric dressing, driven couplers, or flux-tunable bridges, also underscore the landscape of possible interaction-engineering tools to enhance cross-Kerr strength [4, 27, 29, 68, 39]. In general, these approaches introduce additional control complexity, Kerr drifts, or flux-sensitive dissipation, all of which need carefully engineered to reconcile with a high-$Q$ 3D cavity. Therefore, the passive two-qubit bridge may offer a stable, calibration-light, and hybrid cQED-compatible path to achieving strong cross-Kerr without compromising cavity coherence or introducing additional noise channels.
III.2.2 Impact of Qubit-Mode Couplings on Visibility
Figure 5: Qubit coherence visibility $\mathcal{V}_{\mathrm{R}}(\tau_{R})$ as a function of the Ramsey mapping interval $\tau_{R}$ for different dispersive coupling strengths. (a) Dependence on the probe-mode dispersive coupling $\chi_{b}/2\pi=10,\,20,\,50~\mathrm{kHz}$. Larger $\chi_{b}$ accelerates photon-number-induced dephasing during the mapping and therefore shortens the usable Ramsey mapping window (solid curves). A Hahn-echo sequence (dashed curves) suppresses quasi-static photon-number fluctuations and significantly extends the allowable $\tau_{R}$. (b) Dependence on the thermal-mode dispersive coupling $\chi_{a}/2\pi=1,\,5,\,10~\mathrm{kHz}$. Maintaining small $\chi_{a}$ ensures that absorber fluctuations do not appreciably shorten the qubit Ramsey window.
During coherence mapping in the CM scheme, the qubit remains in its ground state throughout the interaction interval $\tau$, and is only promoted to a superposition during the short Ramsey (or Hahn-echo) window $\tau_{R}$ to sample the probe phase. Dephasing in this window originates from photon-number fluctuations of any dispersively coupled mode $m\in\{a,b\}$, with variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$ (Appendix F). The qubit Ramsey coherence envelope reads [25]
$C_{q}=\exp\left\{-4\chi_{m}^{2}\mathrm{Var}(n_{m})\left[\frac{\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}}}{\kappa_{m}^{2}}\right]\right\},$ (9)
where $\mathrm{Var}(n_{b})=|\alpha|^{2}+\bar{n}_{b}(1+\bar{n}_{b})$ for the coherent probe and $\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a})$ for the thermal absorber.
Fig. 5(a) shows that stronger qubit-probe coupling $\chi_{b}$ increases the qubit dephasing rate for a modest probe population $|\alpha|^{2}=1$, yielding a preserved mapping window $\tau_{R}\leq 1~\mu\mathrm{s}$ for the Ramsey envelope (solid lines). When a Hahn-echo sequence is applied, the corresponding filter function, $F_{\mathrm{echo}}(\omega)=8\sin^{4}(\omega\tau_{R}/4)/\omega^{2},$ suppresses the low-frequency photon-number noise that dominates in high-$Q$ probe cavities ($\kappa_{b}\tau_{R}\ll 1$[33, 14, 24]. Therefore, the echo reduces the effective dephasing rate by a factor $(\kappa_{b}\tau_{R})^{2}/32$, restoring near-unity visibility even for $\chi_{b}/2\pi=50~\mathrm{kHz}$ and $\tau_{R}$ can be up to $\sim 6~\mu\mathrm{s}$ (dashed lines). Thus, fast mapping can be performed without appreciable loss of contrast, provided that the probe number fluctuations remain slow compared to $\tau_{R}$.
Fig. 5(b) illustrates that finite qubit-thermal coupling $\chi_{a}$ degrades visibility if present during mapping. Maintaining $\chi_{a}/2\pi\!\lesssim\!1$ kHz ensures that absorber fluctuations do not appreciably shorten the Ramsey window.
By contrast, the same echo operation in the qubit-only scheme would cancel the low-frequency absorber noise that encodes the temperature information, thereby recovering contrast but suppressing thermometric sensitivity.
III.3 Practical considerations and calibration
We next outline general experimental considerations and practical operating procedures that can be used in future implementations to bound the achievable sensitivity and guide calibration. In practice, achievable sensitivity is bounded by visibility loss during the mapping window, incomplete absorber thermalization, excess fluctuations, and systematic parameter errors. The mapping must be completed within the relevant coherence/linewidth scales, which yields a known visibility envelope and an optimal interaction time that maximizes the temperature-dependent slope (Eq. (9) and Fig. 5). Finite absorber equilibration to the engineered bath reduces the available temperature-dependent moments and can be captured by partially thermalized moments set by the absorber-bath coupling and the cycle time (Sec. III.1). Excess noise enters through the same absorber moments and can be modeled as additive offsets, $\bar{n}_{a_{1}}(T)\!\rightarrow\!\bar{n}_{a_{1}}^{\rm th}(T)+\bar{n}_{\rm ex}$ and $\mathrm{Var}(n_{a_{1}})\!\rightarrow\!\mathrm{Var}_{\rm th}(T)+\mathrm{Var}_{\rm ex}$, calibrated via mapping-disabled reference runs (e.g., $\alpha\!=\!0$) and/or fits to known temperature points. Systematic errors (e.g., uncertainty in $\phi=\lambda\tau$ and residual higher-level/weak-drive effects) are mitigated by operating in the dispersive, weak-drive regime and by in situ calibrations of $\phi$ and mode populations (Sec. III.2.1 and App. E/F). These considerations do not alter the operating principle, but they set the optimal working point and provide a straightforward route to separating thermal signal from calibrated offsets.
@@ IV Comparative Performance of Quantum Thermometry Strategies
Figure 6: Comparison of quantum thermometry strategies at fixed cross-Kerr coupling $\lambda/2\pi=50~\mathrm{kHz}$. (a) QFI $\mathcal{F}$ versus sensing time $\tau$ for the CM (blue), PS (orange), and QB (green) schemes. (b) Corresponding coherence envelope visibilities envelope $\mathcal{V}_{\mathrm{coh}}$: probe coherence for CM (blue), constant visibility for PS (orange), and qubit Ramsey visibility for QB (green). Markers indicate the corresponding $\mathcal{V}_{\mathrm{coh}}$ at optimal $\tau^{\star}$ that maximizes $\mathcal{F}$ in CM and QB schemes.
To benchmark performance, we compare the QFI $\mathcal{F}$ and coherence $\mathcal{V}_{\mathrm{coh}}$ across the three schemes as a function of the interaction time $\tau$. Details of the QFI and coherence derivation of the qubit-only (QB) thermometry is presented in Appendix. D. Figure 6 summarizes the results for a fixed cross-Kerr coupling $\lambda/2\pi=50~\mathrm{kHz}$ in the quasi-static regime $\kappa_{a}\tau\ll 1$. This representation isolates the role of coherence decay and highlights the distinct mechanisms that determine each scheme’s optimal operating point.
The QB strategy (green) exhibits the characteristic rise-and-fall behavior expected when the qubit serves as both sensor and memory. For short interaction times, the QFI grows as $\mathcal{F}\propto\tau^{2}$ due to increased phase accumulation. At longer times, however, the same thermal fluctuations that encode temperature also induce qubit dephasing, suppressing the Ramsey visibility and causing $\mathcal{F}$ to collapse. Because this dephasing carries the temperature information itself, echo protocols cannot mitigate it without erasing the signal. The resulting optimal time $\tau^{\star}$ (green marker) defines a narrow high-information window.
The CM strategy (blue) exhibits a qualitatively similar peak-and-rolloff with $\tau$, but with a fundamentally different origin. Here the probe cavity, not the qubit, experiences temperature-induced phase diffusion. During the sensing interval the qubit is fully decoupled, therefore its intrinsic coherence is protected, allowing the probe to absorb all visibility loss. As $\tau$ increases, thermal fluctuations deepen the probe’s phase modulation, boosting $\mathcal{F}$ until probe dephasing dominates and visibility drops. The optimal $\tau^{\star}$ (blue marker) is therefore set by the probe quality factor and operating-point amplitude, both of which are experimentally tunable and ultimately constrained by measurement-induced backaction and finite readout SNR/contrast. This separation between sensing probe and readout qubit enables CM to maintain high per-shot QFI without sacrificing qubit contrast.
The PS tracking scheme (orange) grows monotonically with $\tau$ at the rate expected from deterministic dispersive phase accumulation. Because the measurement does not rely on coherence of either the qubit or the probe, visibility remains essentially unity, and $\mathcal{F}_{T}$ lacks the peak-collapse structure associated with the coherence-based approaches. While its instantaneous QFI is lower at short times, the PS scheme can integrate indefinitely and is therefore limited primarily by measurement efficiency and low-frequency technical drifts.
For CM and QB schemes, the scaling $\mathcal{F}_{T}(\tau)\sim\tau^{2}e^{-2\Gamma\tau}$ dictates an optimal interaction time $\tau^{\star}\sim 1/2\Gamma$, where $\Gamma$ is the relevant dephasing rate (probe dephasing for CM, or qubit dephasing for QB). This defines an effective bandwidth $B^{\star}\sim\Gamma$ for rapid temperature tracking. By contrast, the PS strategy has $\mathcal{F}_{T}(\tau)\sim\tau^{2}$ without a visibility penalty, enabling long integrations but without the coherence-assisted gain available to CM.
The CM thermometry combines the short-time sensitivity of coherence-based sensing with resilience to the self-dephasing bottleneck that limits QB. The probe absorbs the visibility cost, the qubit retains full coherence contrast, and the QFI remains high across a broad temporal window. QB provides comparable sensitivity at longer times subsequently experiences rapid visibility collapse. PS is the most robust for long integrations but lacks the coherence-amplification available to CM. These results collectively position CM as the most versatile operating mode for quantum-limited thermometry, offering strong per-shot information and robustness to over-coupling and parameter drift.